Session: Feb 2005 (session 1)
Wavelength: 18 cm
To give feedback on this experiment, click
Urumqi (Liu Xiang at Wed Mar 2 02:25:31 2005) Minor failures : TSYS
Torun (K.Borkowski at Thu Mar 3 07:53:04 2005) Minor failures :
Cal data gathered from 3.5 hour observations exhibit large
scatter. Probably we will have to repeat these measurements.
Medicina (Andrea Orlati at Thu Mar 3 09:03:22 2005) Success :
Hartebeesthoek (Jonathan Quick at Thu Mar 3 12:19:26 2005) Minor failures : RFI
Some reasonable data despite RFI, RXG file updated.
Onsala (Michael Lindqvist at Fri Mar 4 08:40:48 2005) Success :
Cooling problem probably solved. Data obtained both for warm and cooled HEMTs.
Jodrell_Bank (Alastair Gunn at Mon Mar 7 11:59:56 2005) Success :
The Lovell Telescope was used.
Good data obtained for 18cm band only. Will try to get 21cm as well.
Noto (P. Cassaro at Mon Mar 7 15:30:58 2005) Success :
Effelsberg (A. Kraus at Fri Mar 11 12:29:40 2005) Success :